# Multiple Instance Learning for Glioma Diagnosis using Hematoxylin and Eosin stained Whole Slide Images: An Indian Cohort Study

### 24 Feb 2024 v1

@@ Abstract
Brain tumors represent a severe and life-threatening condition, demanding precise diagnosis and tailored treatment strategies. This study advances patient care with findings from rigorous multiple-instance-learning experimentations across various feature extractors, and aggregators in brain tumor histopathology. It establishes new performance benchmarks in glioma subtype classification across multiple datasets, including a novel dataset focused on the Indian demographic (IPD-Brain), providing a valuable resource for existing research. Using a ResNet-50, pretrained on histopathology datasets, for feature extraction, combined with DTFD feature aggregator, our approach achieves state-of-the-art AUCs of 88.08 ± 3.98 on IPD-Brain and 95.81 ± 1.78 on TCGA-Brain dataset respectively for three-way glioma subtype classification. Moreover, it establishes new benchmarks in grading and detecting IHC molecular biomarkers (IDH1 (mutant R132H), TP53, ATRX, Ki-67) through H&E stained whole slide images for the IPD-Brain dataset. The work also highlights a significant correlation between the model’s decision-making processes and the diagnostic reasoning of pathologists, underscoring its capability to mimic professional diagnostic procedures.
@@ 1 Introduction
Brain tumors, known for their lethality, represent a critical medical condition characterized by the abnormal growth of cells within the central nervous system, which can lead to severe neurological dysfunction. According to the World Health Organisation (WHO), central nervous system (CNS) tumors have extensive classification and subtypes. Due to genetic predisposition, environmental factors, and lifestyle-related variables in India, the occurrence of CNS tumor cases fluctuates between 5 and 10 cases per 100,000 individuals [1, 2]. Glial tumors are the most common type of brain tumor, primarily Astrocytoma, Glioblastoma, and Oligodendroglioma [3, 4, 5]. Glioblastoma, the most aggressive among them, which accounts for approximately 57% of all gliomas, represents a highly malignant glioma, with a median survival of just 12-15 months even with aggressive treatment approaches and associated with a high rate of recurrence due to its infiltrative nature [6, 7]. Astrocytoma, on the other hand, originates from astrocytes and is graded based on its degree of malignancy, low or high. Oligodendroglioma, a rarer subtype, arises from oligodendrocytes and is often characterized by genetic alterations, which can influence treatment strategies and predict better outcomes [8].
@@
Figure 1: Multi-Instance Learning Framework for Subtype Classification in Brain Histopathology. Tissue from surgical samples is digitized and processed into patches, and patch-level features are extracted using a pre-trained feature extractor network. Subsequently, the feature aggregation method pools the patch-level representation into slide-level representation, which is then classified to determine tumor type, grade, and molecular markers. Then attention mapping is used for interpretability, providing critical insights into the neural network’s decision-making process.
@@
Additionally, IHC staining is a specialized technique to detect specific proteins or antigens within tissue samples. It involves using antibodies tagged with chromogenic or fluorescent markers to visualize the presence and distribution of specific proteins in tissues. IHC is valuable for investigating biomarkers like isocitrate dehydrogenase (IDH / IDH1 (mutant R132H)), alpha thalassemia/mental retardation syndrome X-linked (ATRX), TP53, Ki-67 and identifying specific cellular components related to diseases. In this study, we explored the mutation status of IDH, ATRX, TP53, and Ki-67 using H&E stain to significantly alleviate the workload in pathology labs and reduce associated costs. By diminishing the need for multiple IHC tests, this approach conserves resources, removes the necessity since it may not be available in all healthcare systems, and streamlines the diagnostic process. To mimic the diagnosis of pathologists, we treated Ki-67, the percentage of positively stained cells among the total number of cancer cells assessed, as a binary classification task by varying thresholds.
@@
In recent years, the field of histopathology has undergone a significant transformation, primarily due to the emergence of digital histopathology [14], the available datasets like TCGA [15], and the application of deep learning methods for their analysis [16, 17]. This evolution in the brain tumor research domain underscores the importance of diverse datasets that extend beyond the confines of American or European populations. The global landscape of brain tumors demands a comprehensive understanding facilitated by including datasets from various demographics. Deep learning methods have demonstrated remarkable efficiency in processing extensive histopathological data with unprecedented speed and accuracy.
@@
While existing research is able to predict the subtype, grading, and biomarkers directly from routine H&E stained histopathology slides for many tumor types [18, 19, 16], including brain tumor [20, 21], showing the potential to accelerate diagnostic workflows and reduce the costs associated with molecular testing. However, a limited focus is still given to brain tumor classification based on the updated WHO guidelines (the fifth edition, WHO CNS5, 2021) [22]. Due to the large size of histopathology images and lack of patch-level labels in a whole slide image (WSI), the Multiple Instance Learning (MIL) framework has been widely adopted in the field of digital pathology. MIL is a two-stage framework: (i) Feature extractor and (ii) Feature Aggregator. The use of standard ImageNet pre-trained weights and large-scale domain-aligned (pathology) pre-training in a self-supervised manner [23, 24, 25] for extracting latent features of patches underscores the effectiveness on various downstream pathology tasks. To aggregate or pool these features, Campanella et al. [18] employed an RNN-based approach for feature-level aggregation in the classification of clinical pathology whole slide images. ABMIL, as introduced by Ilse et al. [19], leverages attention mechanisms to determine instance weights through a trainable neural network. This method enhances traditional MIL pooling operations, such as max or mean pooling, by aggregating bag-level representations based on weighted averages of instance-level features. Expanding on the concept of attention-based instance scoring, CLAM, developed by Lu et al. [16], delves into multi-class predictions using instance-level clustering, thereby facilitating the learning of class-specific features. Dual-stream MIL (DS-MIL) [26] employs non-local attention for aggregation to learn classifiers. A Transformer-based Multiple Instance Learning approach (TransMIL) [27] explores the relationships between instances in terms of their morphology and spatial characteristics using the Nyström method for approximating all-to-all self-attention, though it requires significant computation resources. DTFD-MIL, introduced by Zhang et al. [28], introduces the concept of ”pseudo-bags” addressing issues associated with small cohorts. Another contribution is a transformer-based hierarchical ViT proposed by [25], focusing on using spatial relationships among instances. However, despite these advancements, the downstream classification tasks involving brain tumor subtypes and the identification of immunohistochemistry (IHC) biomarkers remain underexplored.
@@
Our research addresses a critical gap in brain tumor studies by introducing a glioma subtype dataset within the India Pathology Dataset (IPD) consortium. [^1] https://hai.iiit.ac.in/ipd/ This dataset facilitates research in India and encourages global collaboration to advance our understanding and combat this lethal affliction. Fig. 1 shows the pipeline followed in this work. Through rigorous experimentation, we aim to identify a specific combination of feature extractor and aggregator that performs well for various tasks based on the 2021 WHO guidelines. The outcomes of our study emphasize the opportunity to advance the state of the art in deep learning applications for brain tumor classification.
@@
  • We created a dataset tailored to the Indian region and one of the largest collections of brain whole slide images globally, providing a valuable resource for research in this domain.
  • To our knowledge, this is the first study predicting Immunohistochemistry (IHC) molecular biomarkers, utilizing cost-effective Haematoxylin and Eosin (H&E) slides, eliminating the need for expensive IHC staining. This not only enhances accessibility but also explores the potential of H&E stains in accurately determining these biomarkers.
  • We achieve state-of-the-art (SOTA) results on The Cancer Genome Atlas (TCGA) data for glioma subtype classification by using a specific self-supervised pre-trained feature extractor and a feature aggregator.
@@ 3.1 MIL Problem formulation
In histopathology classification, Multiple Instance Learning (MIL) is a compelling solution to the challenge of sparse and imprecise annotations often encountered in medical image analysis. Unlike traditional supervised learning, which relies on precisely labeled instances (patches), MIL operates on ’bags’ of instances, with the label determined by the collective characteristics of these instances (slide-level). Here, the mathematical formulation of the MIL framework for subtype classification is extended and defined as follows:
Consider a dataset $X$ comprising $M$ WSIs, denoted as $X=\{X_{1},X_{2},...,X_{M}\}$. Each WSI is a ’bag’ in MIL terminology, representing an entire histopathology slide. The label of the $m$-th WSI, denoted as $Y_{m}$, in an n-way subtype classification task, is determined based on the subtype presence within the slide, with each slide categorized into one of $n$ possible subtypes or negative slide. The instances within a bag denoted as $x_{i}$, are individual image patches or regions extracted from the histopathology slide. For a given $m$-th WSI, the instances are represented below in (1):
@@ 3.2 WSI Preprocessing
The process begins by segmenting tissue regions within digitized slides. The image is converted to the HSV color space, and a mask is created by thresholding the saturation channel after smoothing. Morphological operations fill gaps and holes, and contours are filtered based on area. We have included optimal segmentation parameters for our dataset in the section 4. Following segmentation, the algorithm crops 256x256 patches from within the segmented regions based on user-specified magnification.
@@ 3.5 Transfer Learning for IHC Biomarkers
Furthermore, our investigation aimed to assess ”Is H&E stain enough?” by determining the capability of our weakly-supervised deep learning subtype classification model in predicting common diagnostic molecular markers from glioma H&E slides. We sought to evaluate its potential utility as a clinical secondary check or a cost-effective alternative to actual molecular biology experiments. Using transfer learning, a technique that adapts a model trained on one task to a related new task, we systematically explored the effectiveness of the best-performing model, initially trained for tumor subtype classification, in inferring molecular biomarkers within a weakly supervised context. Fine-tuning was conducted on the three-class subtyping model for each of three binary classifications of IHC mutations (somatic mutation versus wild type for IDH1, TP53, and ATRX mutations) and one classification task at various cutoffs for assessing tumor cell proliferation (Ki-67).
@@ 4.1 Experiment Setup
In our study, we conducted comprehensive experiments using two datasets: the in-house IPD-Brain Dataset and the publically available TCGA-Brain. This section delineates the experimental setup employed. For feature extraction at the patch level from WSIs, we used four distinct networks: ResNet-50 pre-trained on ImageNet in a supervised manner, ResNet-50 pre-trained using the Barlow Twins self-supervised method for histopathology, CTransPath pre-trained with semantically relevant contrastive learning for histopathology, and ViT-S pre-trained via the DINO approach for histopathology.
Upon extracting features using these models, we applied three aggregation methods - CLAM, DSMIL, and DTFD - for the downstream classification task. We identified the most effective extractor-aggregator combination on the IPD-Brain dataset for the three-way subtype classification task and subsequently extended its application for further experimental analysis. This involved first testing the chosen combination on the TCGA-Brain dataset for subtype classification, followed by employing the same combination for classifying grading (LGG vs HGG) and IHC biomarkers mutations (TP53, IDH, ATRX, Ki-67) using H&E stained slides. This systematic approach allowed us to rigorously assess the efficacy of our models across different datasets and tasks for brain histopathology.
@@ 5 Results
Our analysis across various datasets demonstrates the effectiveness of combining feature extractors, aggregators, and pretraining techniques. On the India Pathology Dataset Brain (IPD-Brain), Table 2, the combination of a Resnet-50 backbone, pretrained via the Barlow twin self-supervised learning (SSL) method, and the DTFD aggregator emerged as superior, achieving an average AUC of 88.08%, an accuracy of 79.83%, and an F1 score of 75.87%. Variation in AUC and F1 score across 10-folds for all the combinations can be observed in Fig. 4. This configuration proved particularly adept at classifying brain tumor subtypes, showcasing its suitability for the unique challenges of histopathological data analysis.
@@
Table 5 details the classification of Immunohistochemistry (IHC) biomarkers on the IPD-Brain dataset, achieving high accuracy and AUC rates for crucial biomarkers like IDH, ATRX, and TP53. This highlights the model’s potential as an accessible IHC molecular biomarker identification tool directly from H&E slides.
@@
Furthermore, Table 6, the Ki-67 marker classification task showcased the model’s capability to accurately predict tumor cell proliferation rates, with significant proficiency observed at the cutoff value of 10, achieving an 89.55% AUC and a 91.17% F1 score. This proficiency in Ki-67 classification underlines the model’s utility in providing vital prognostic information necessary for effective treatment planning and patient management.
@@
Our findings demonstrate that a ResNet-50 model pre-trained with Barlow Twin SSL and combined with DTFD aggregation enhances histopathological analysis, especially in brain tumor classification. This approach underscores the effectiveness of careful feature extraction method selection, showing that proper pretraining and aggregation negate the need for stain normalization and improve diagnostic accuracy.
@@ 6 Explainability
Our model’s practical applicability and accuracy were further substantiated through a rigorous single-blind evaluation conducted with a seasoned pathologist from the Department of Pathology at Nizam’s Institute of Medical Sciences, a premier tertiary hospital in India. The pathologist was presented with a WSI as our model. The pathologist scrutinized specific regions within the WSIs to formulate their diagnoses, some of which are shown as ’Raw Patch’ in Fig. 5.
@@
By qualitatively analyzing, we can observe the DTFD’s superior ability to focus on regions indicative of each glioma’s medical characteristics, aligning with known pathological features. For Glioblastoma, it highlights areas showing its aggressive nature, like microvascular proliferation and pseudopalisading necrosis. In Astrocytoma, the emphasis is on cellular structures absent of necrosis, showcasing gemistocytic features. Oligodendroglioma’s distinctive ’fried egg’ appearance and ’chicken wire’ vascular pattern are accurately identified, distinguishing it from other subtypes. This assessment underlines DTFD’s effectiveness in pinpointing diagnostically relevant regions, enhancing model transparency and reliability in histopathological evaluations.
The model demonstrated proficiency in identifying critical morphological microstructures, showing a strong alignment between its focus on important regions and the key areas within the raw whole slide images (WSIs) that are essential for glioma subtype classification. This congruence closely reflects the diagnostic strategies employed by pathologists, underscoring the model’s ability to prioritize and analyze patch segments in a manner similar to how experts identify glioma patterns.

### 08 Mar 2024 v2

# Multiple Instance Learning for Glioma Diagnosis using Hematoxylin and Eosin Whole Slide Images: An Indian Cohort Study

@@ Abstract
The effective management of brain tumors relies on precise typing, subtyping, and grading. This study advances patient care with findings from rigorous multiple-instance-learning experimentations across various feature extractors and aggregators in brain tumor histopathology. It establishes new performance benchmarks in glioma subtype classification across multiple datasets, including a novel dataset focused on the Indian demographic (IPD-Brain), providing a valuable resource for existing research. Using a ResNet-50, pretrained on histopathology datasets for feature extraction, combined with the Double-Tier Feature Distillation (DTFD) feature aggregator, our approach achieves state-of-the-art AUCs of 88.08 $\pm$ 3.98 on IPD-Brain and 95.81 $\pm$ 1.78 on the TCGA-Brain dataset, respectively, for three-way glioma subtype classification. Moreover, it establishes new benchmarks in grading and detecting IHC molecular biomarkers (IDH1R132H, TP53, ATRX, Ki-67) through H&E stained whole slide images for the IPD-Brain dataset. The work also highlights a significant correlation between the model’s decision-making processes and the diagnostic reasoning of pathologists, underscoring its capability to mimic professional diagnostic procedures.
@@ 1 Introduction
Brain tumors, known for their lethality, represent a critical medical condition characterized by the abnormal growth of cells within the central nervous system (CNS), which can lead to severe neurological dysfunction. The World Health Organization (WHO) Classification of CNS tumors (WHO CNS 5) includes various types and subtypes based on their histologic and molecular genetic characteristics [9]. Due to genetic predisposition, environmental factors, and lifestyle-related variables in India, the occurrence of CNS tumor cases fluctuates between 5 and 10 cases per 100,000 individuals [1, 2]. Glial tumors are the most common type of brain tumor, primarily Astrocytoma, Glioblastoma, and Oligodendroglioma [3, 4, 5]. Glioblastoma, the most aggressive among them, which accounts for approximately 57% of all gliomas, represents a highly malignant glioma, with a median survival of just 12-15 months, even with aggressive treatment approaches and associated with a high rate of recurrence due to its infiltrative nature [6, 7]. Astrocytoma, on the other hand, originates from astrocytes and is graded based on its degree of malignancy, low or high. Oligodendroglioma, a rarer subtype, arises from oligodendrocytes and is often characterized by genetic alterations, which can influence treatment strategies and predict better outcomes [8].
@@
Figure 1: Multi-Instance Learning Framework for Subtype Classification in Brain Histopathology. Tissue from surgical samples is digitized and processed into patches, and patch-level features are extracted using a pre-trained feature extractor network. Subsequently, the feature aggregation method pools the patch-level representation into slide-level representation, which is then classified to determine tumor type, grade, and molecular markers. Then, attention mapping is used for interpretability, providing critical insights into the neural network’s decision-making process.
@@
Additionally, IHC staining is a specialized technique to detect specific proteins or antigens within tissue samples. It involves using antibodies tagged with chromogenic or fluorescent markers to visualize the presence and distribution of specific proteins in tissues. IHC is valuable for investigating biomarkers like isocitrate dehydrogenase (IDH / IDH1R132H), alpha thalassemia/mental retardation syndrome X-linked (ATRX), TP53, Ki-67 and identifying specific cellular components related to diseases. In this study, we explored the possibility of predicting the status of IHC biomarkers (IDH, ATRX, TP53, and Ki-67) using H&E stained whole slide images (WSIs). This will help to understand between morphology and genetics of the tumor and perhaps help in assisting the diagnosis and reducing multiple molecular assessments using IHC.
@@
In recent years, the field of histopathology has undergone a significant transformation, primarily due to the emergence of digital histopathology [14], the available datasets like The Cancer Genome Atlas (TCGA) [15], and the application of deep learning methods for their analysis [16, 17]. This evolution in brain tumor research domain underscores the importance of diverse datasets that extend beyond the confines of American or European populations. The global landscape of brain tumors demands a comprehensive understanding facilitated by including datasets from various demographics. Deep learning methods have demonstrated remarkable efficiency in processing extensive histopathological data with unprecedented speed and accuracy.
@@
While existing research is able to predict the subtype, grading, and biomarkers directly from routine H&E stained histopathology slides for many tumor types [18, 19, 16], including brain tumor [20, 21], showing the potential to accelerate diagnostic workflows and reduce the costs associated with molecular testing. However, a limited focus is still given to brain tumor classification based on the updated WHO guidelines (the fifth edition, WHO CNS5, 2021) [22]. Due to the large size of histopathology images and lack of patch-level labels in a WSI, the Multiple Instance Learning (MIL) framework has been widely adopted in the field of digital pathology. MIL is a two-stage framework: (i) Feature extractor and (ii) Feature Aggregator. The use of standard ImageNet pre-trained weights and large-scale domain-aligned (pathology) pre-training via self-supervised learning (SSL) [23, 24, 25] for extracting latent features of patches underscores its effectiveness on various downstream pathology tasks. To aggregate or pool these features, Campanella et al. [18] employed an RNN-based approach for feature-level aggregation in the classification of clinical pathology whole slide images. ABMIL, as introduced by Ilse et al. [19], leverages attention mechanisms to determine instance weights through a trainable neural network. This method enhances traditional MIL pooling operations, such as max or mean pooling, by aggregating bag-level representations based on weighted averages of instance-level features. Expanding on the concept of attention-based instance scoring, Clustering-constrained Attention MIL (CLAM), developed by Lu et al. [16], delves into multi-class predictions using instance-level clustering, thereby facilitating the learning of class-specific features. Dual-stream MIL (DS-MIL) [26] employs non-local attention for aggregation to learn classifiers. A Transformer-based Multiple Instance Learning approach (TransMIL) [27] explores the relationships between instances in terms of their morphology and spatial characteristics using the Nyström method for approximating all-to-all self-attention, though it requires significant computation resources. Double-Tier Feature Distillation (DTFD) MIL, introduced by Zhang et al. [28], introduces the concept of ”pseudo-bags” addressing issues associated with small cohorts. Another contribution is a transformer-based hierarchical ViT proposed by [25], focusing on using spatial relationships among instances. However, despite these advancements, the downstream classification tasks involving brain tumor subtypes and the identification of IHC biomarkers remain underexplored.
@@
Our research addresses a critical gap in brain tumor studies by introducing a glioma subtype dataset within the India Pathology Dataset (IPD) consortium. [^1] https://hai.iiit.ac.in/ipd/ This dataset facilitates research in India and encourages global collaboration to advance our understanding and combat this severe illness. Fig. 1 shows the pipeline followed in this work. Through rigorous experimentation, we aim to identify a specific combination of feature extractor and aggregator that performs well for various tasks based on the WHO CNS 5 guidelines. The outcomes of our study emphasize the opportunity to advance the state of the art in deep learning applications for brain tumor classification.
@@
  • We created a dataset tailored to the Indian region and one of the largest collections of brain WSIs globally, providing a valuable resource for research in this domain.
  • To our knowledge, this is the first study predicting IHC molecular biomarkers using cost-effective H&E slides, eliminating the need for expensive IHC staining. This enhances accessibility and also explores the potential of H&E stains in accurately determining these biomarkers.
  • We achieve state-of-the-art (SOTA) results on TCGA data for glioma subtype classification by using a specific self-supervised pre-trained feature extractor and a feature aggregator.
@@
  • We introduce Ki-67, the percentage of positively stained cells among the total number of cancer cells assessed, as a binary classification task by varying thresholds.
@@ 3.1 MIL Problem formulation
In histopathology classification, MIL is a compelling solution to the challenge of sparse and imprecise annotations often encountered in medical image analysis. Unlike traditional supervised learning, which relies on precisely labeled instances (patches), MIL operates on ’bags’ of instances, with the label determined by the collective characteristics of these instances (slide-level). Here, the mathematical formulation of the MIL framework for subtype classification is extended and defined as follows:
Consider a dataset $X$ comprising $M$ WSIs, denoted as $X=\{X_{1},X_{2},...,X_{M}\}$. Each WSI is a ’bag’ in MIL terminology, representing an entire histopathology slide. The label of the $m$-th WSI, denoted as $Y_{m}$, in an n-way subtype classification task, is determined based on the subtype presence within the slide, with each slide categorized into one of $n$ possible subtypes or negative slide. The instances within a bag, denoted as $x_{i}$, are individual image patches or regions extracted from the histopathology slide. For a given $m$-th WSI, the instances are represented below in (1):
@@ 3.2 WSI Preprocessing
The process begins by segmenting tissue regions within digitized slides. The image is converted to the HSV color space, and a mask is created by thresholding the saturation channel after smoothing. Morphological operations fill gaps and holes, and contours are filtered based on area. We have included optimal segmentation parameters for our dataset in section 4. Following segmentation, the algorithm crops 256x256 patches from within the segmented regions based on user-specified magnification.
@@ 3.5 Transfer Learning for IHC Biomarkers
Furthermore, our investigation aimed to assess ”Is H&E stained slide enough?” by determining the capability of our weakly-supervised deep learning subtype classification model in predicting common diagnostic molecular markers from glioma H&E slides. We sought to evaluate its potential utility as a clinical secondary check or a cost-effective alternative to actual molecular biology experiments. Using transfer learning, a technique that adapts a model trained on one task to a related new task, we systematically explored the effectiveness of the best-performing model, initially trained for tumor subtype classification, in inferring molecular biomarkers within a weakly supervised context. Fine-tuning was conducted on the three-class subtyping model for each of the three binary classifications of IHC mutations (somatic mutation versus wild type for IDH1, TP53, and ATRX mutations) and one classification task at various cutoffs for assessing tumor cell proliferation (Ki-67).
@@ 4.1 Experiment Setup
In our study, we conducted comprehensive experiments using two datasets: the in-house IPD-Brain Dataset and the publically available TCGA-Brain. This section delineates the experimental setup employed. For feature extraction at the patch level from WSIs, we used four distinct networks: ResNet-50 pre-trained on ImageNet in a supervised manner, ResNet-50 pre-trained using the Barlow Twins SSL method for histopathology, CTransPath pre-trained with semantically relevant contrastive learning for histopathology, and ViT-S pre-trained via the DINO approach for histopathology.
Upon extracting features using these models, we applied three aggregation methods - CLAM, DSMIL, and DTFD - for the downstream classification task. We identified the most effective extractor-aggregator combination on the IPD-Brain dataset for the three-way subtype classification task and subsequently extended its application for further experimental analysis. This involved first testing the chosen combination on the TCGA-Brain dataset for subtype classification, followed by employing the same combination for classifying grading (LGG vs HGG) and IHC biomarkers mutations (TP53, IDH, ATRX, Ki-67) using H&E stained slides. This systematic approach allowed us to assess the efficacy of our models rigorously across different datasets and tasks for brain histopathology.
@@ 5 Results
Our analysis across various datasets demonstrates the effectiveness of combining feature extractors, aggregators, and pretraining techniques. On the India Pathology Dataset Brain (IPD-Brain), Table 2, the combination of a Resnet-50 backbone, pretrained via the Barlow twin SSL method, and the DTFD aggregator emerged as superior, achieving an average AUC of 88.08%, an accuracy of 79.83%, and an F1 score of 75.87%. Variation in AUC and F1 score across 10-folds for all the combinations can be observed in Fig. 4. This configuration proved particularly adept at classifying brain tumor subtypes, showcasing its suitability for the unique challenges of histopathological data analysis. For the grade classification (LGG vs HGG) task on IPD-Brain, the same configuration achieved an AUC of 92.64%, accuracy of 89.25%, and F1 of 92.79%.
@@
Table 5 details the classification of IHC molecular biomarkers on the IPD-Brain dataset, achieving high accuracy and AUC rates for crucial biomarkers like IDH, ATRX, and TP53. This highlights the model’s potential as an accessible IHC molecular biomarker identification tool directly from H&E slides.
@@
Furthermore, in Table 6, the Ki-67 marker classification task showcased the model’s capability to accurately predict tumor cell proliferation rates, with significant proficiency, observed at the cutoff value of 10, achieving an 89.55% AUC and a 91.17% F1 score. This proficiency in Ki-67 classification underlines the model’s utility in providing vital prognostic information necessary for effective treatment planning and patient management.
@@
Our findings demonstrate that a ResNet-50 model pre-trained with Barlow Twin SSL and combined with DTFD aggregation enhances morphological and immunohistochemical analysis in histopathology, especially in brain tumor classification. This approach underscores the effectiveness of careful feature extraction method selection, showing that proper pretraining and aggregation negate the need for stain normalization and improves diagnostic accuracy.
@@ 6 Explainability
Our model’s practical applicability and accuracy were further substantiated through a rigorous single-blind evaluation conducted with an experienced neuropathologist from the Department of Pathology at Nizam’s Institute of Medical Sciences, a premier tertiary hospital in India. The pathologist was presented with a WSI as our model. The pathologist scrutinized specific regions within the WSIs to formulate their diagnoses, some of which are shown as ’Raw Patch’ in Fig. 5.
@@
By qualitatively analyzing, we can observe DTFD’s superior ability to focus on regions indicative of each glioma’s medical characteristics, aligning with known pathological features. The key diagnostic features of glioblastoma, such as microvascular proliferation and palisading necrosis, were accurately highlighted by DTFD. In astrocytoma, the emphasis is on cellular structures absent of necrosis, showcasing gemistocytic features. The classic histopathologic features of oligodendroglioma, like the ”chicken wire” vascular pattern and ”fried egg” appearance of the cells, are also accurately identified. This assessment underlines DTFD’s effectiveness in pinpointing diagnostically relevant regions, enhancing model transparency and reliability in histopathological evaluations.
The model demonstrated proficiency in identifying critical morphological microstructures, showing a strong alignment between its focus on important regions and the key areas within the raw WSIs that are essential for glioma subtype classification. This congruence closely reflects the diagnostic strategies employed by pathologists, underscoring the model’s ability to prioritize and analyze patch segments in a manner similar to how experts identify glioma patterns.